# GOProteinGNN: Leveraging Protein Knowledge Graphs for Protein Representation Learning

### v1

@@ 4.2.1. Contact Prediction
4.2.2. Remote Homology Detection
@@
4.2.3. Protein-Protein Interaction Identification
Protein-protein interactions (PPI) refer to the physical contacts between two or more amino acid sequences, or two proteins, in our study. The goal is to predict the interaction type(s) of each protein pair, out of 7 possible interactions (i.e., multilabel classification task). The evaluation is conducted on two datasets: SHS27K (Chen et al., 2019) and SHS148K (Chen et al., 2019). Breadth-first search (BFS) and depth-first search (DFS) are used to generate test sets from the two datasets. F1 score is used as the evaluation metric.
4.2.4. Semantic Similarity Inference
This analysis aims to evaluate the extent to which representation models capture information regarding biomolecular functional similarity, with a specific emphasis on the biological process (BP) category within the Gene Ontology (GO) dataset. We adopt the semantic similarity measures outlined by Unsal et al. and compute Spearman’s rank correlation between representation vector similarities and the actual GO-based semantic similarities of corresponding protein pairs. Higher correlation values indicate greater success in representing functional similarity.
4.2.5. Protein–Protein Binding Affinity Estimation
@@
This task involves assessing the ability of a protein representation to predict alterations in binding affinity caused by protein mutations. Each protein pair is associated with a real value, making it a regression task that reflects binding affinity changes. We follow the approach of Unsal et al. and use the SKEMPI dataset (Moal and Fernández-Recio, 2012) with an evaluation based on the mean squared error obtained by 10-fold cross-validation.
4.3. Baselines
Incorporating SOTA models as our baselines, we used ProtBert (Elnaggar et al., 2022) and ESM-1b (Rives et al., 2019) as benchmarks, following KeAP (Zhou et al., 2023) for fair comparison. The former is based on a BERT architecture whereas the latter is based on the AlphaFold (Jumper et al., 2021) architecture adapted for the tasks mentioned in Section 4.2. These models are considered as the top-performing models in the absence of factual knowledge integration. Furthermore, we subjected our analysis to the scrutiny of the most recent and advanced knowledge-driven pre-training model, KeAP (Zhou et al., 2023). We draw the reader’s attention that we exclude the results of OntoProtein (Zhang et al., 2022) in this paper, as it was presented by KeAP (Zhou et al., 2023) that it is significantly inferior on a dataset without leakage. For all baseline models, we generate representations during the training phase and subsequently assess their performance in various biological tasks. For each specific task, we followed the convention of selecting SOTA methods, just as prior research has done.
4.3.1. Task Specific Baselines
Building upon the KeAP (Zhou et al., 2023) baselines, we extend our evaluation to encompass additional SOTA models that are known to perform well on specific tasks within the protein domain. For the contact and homology tasks we include variants of LSTM (Hochreiter and Schmidhuber, 1997), ResNet (He et al., 2015), and Transformer (Vaswani et al., 2017), as well as ProteinBert (Brandes et al., 2021) for the homology. Following Zhang et al., we examine DPPI (Hashemifar et al., 2018), DNNPPI (Li et al., 2018), PIPR (Chen et al., 2019), and GNN-PPI (Lv et al., 2021) for the protein-protein interactions (PPI) task. Also, we incorporate PIPR (Chen et al., 2019) for affinity estimation and MSA Transformer (Rao et al., 2021) and ProtT5-XL (Elnaggar et al., 2022) for semantic similarity inference.
4.4. Implementation Details
During pre-training, GOProteinGNN is trained for 50k steps using a learning rate of 1e-4, weight decay of 0.02, and a batch size of 512 (proteins) over four GPUs (NVIDIA A40, 48GB Memory each). RGCN layers $L$ are set to 1 and GNN Knowledge Injection (GKI) components (referred to as $k$) are set to 3. For downstream tasks, we use hyperparameters from GNN-PPI (Lv et al., 2021) for PPI prediction and PROBE (Unsal et al., 2022) for binding affinity and similarity inference. For TAPE (Rao et al., 2019), we follow KeAP (Zhou et al., 2023) and perform a grid search to optimize batch size and learning rate. In terms of computational complexity, GOProteinGNN, like other GCNs, operates at approximately $O(|V|*|E|)$. Several optimizations, including limiting the number of hops per protein, reduce the complexity to $O(|V|)$, achieving performance akin to KeAP. During application, the complexity of both algorithms is $O(1)$. For full implementation details, refer to the provided code.
5. Empirical Results
In this section, we present our empirical results. We validate the statistical significance of the results using a one-tailed paired Student’s t-test with a 95% confidence level. Significant results are marked with an asterisk (*) in the tables.
5.1. Contact Prediction
In Table 1, we present GOProteinGNN’s performance on the contact prediction task compared to six baseline methods. GOProteinGNN significantly outperformed all baseline methods across all categories and metrics. In particular, GOProteinGNN showed superior performance compared to ProtBert (Elnaggar et al., 2022) and ESM-1b (Rives et al., 2019), which lack knowledge graph information. This highlights the importance of incorporating knowledge during representation learning, especially for pairwise amino acid tasks (e.g., contact prediction task), where success depends on understanding the proteins’ functions.
Moreover, KeAP (Zhou et al., 2023) incorporates a knowledge post-encoder, limiting its capacity to seamlessly integrate external knowledge into protein representations, as reflected in the performance gains of GOProteinGNN over KeAP. These findings highlight the significant impact of incorporating knowledge graph information in GOProteinGNN via the GKI layers, leading to superior performance in the contact prediction task.
5.2. Remote Homology Detection
In Table 2, we present GOProteinGNN’s performance compared to seven baseline methods on the remote homology detection task.
Despite the challenging nature of the task, due to the large number of potential protein folds, GOProteinGNN significantly outperformed all methods, demonstrating its effectiveness in handling the complexities of remote homology detection. Specifically, GOProteinGNN achieved an advantage over KeAP (Zhou et al., 2023), further affirming its ability to capture essential biological contexts and interactions which is crucial for predicting the protein fold structure, as it processed the whole protein knowledge graph.
As remote homology detection is a sequence-level classification task, we can see that both KeAP (Zhou et al., 2023) and ProtBert (Elnaggar et al., 2022) achieved equal results, suggesting KeAP fails to leverage the factual knowledge in the protein level in this task.
This task highlights the novelty of our algorithm which allows for the integration of information at both the individual amino acid level and the entire protein level.
Table 2. Results on remote homology detection task. The best result is highlighted in bold.
Method Homology
LSTM 0.26
ResNet 0.17
Transformer 0.21
ProtBert 0.29
ProteinBert 0.22
ESM-1b 0.11
KeAP 0.29
GOProteinGNN $\textbf{0.30}^{*}$
5.3. Protein-Protein Interaction Identification
Table 3 presents the results of the protein-protein interaction (PPI) identification task. GOProteinGNN showed strong performance across the board, achieving the highest F1 scores across the majority of datasets and settings with statistical significance, which can be beneficial for drug discovery or personalized medicine.
In particular, for SHS27K, GOProteinGNN obtained SOTA results with an increase of approximately 1% using BFS and 4% using DFS compared to KeAP (Zhou et al., 2023). For the SHS148K dataset (Chen et al., 2019), GOProteinGNN demonstrated superiority over all other models when using BFS, while maintaining its performance as the second-best performer when using DFS. The trend of declining improvement can be attributed to the increasing amount of fine-tuning data (from SHS27K to SHS148K) that reduces the impact of pre-training and the knowledge graph injection.
In conclusion, the results highlight the strengths of our GOProteinGNN, which enable better biological understanding of the protein, leading to success in predicting interactions between proteins.
Table 3. Results on protein-protein interaction (PPI) identification task. The table displays the results in two datasets: SHS27K and SHS148K. BFS and DFS denote the outcomes of breadth-first search and depth-first search methods, respectively. The F1 score is the evaluation metric. The best result is highlighted in bold.
Method SHS27K SHS148K
BFS DFS BFS DFS
DNN-PPI 48.09 54.34 57.40 58.42
DPPI 41.43 46.12 52.12 52.03
PIPR 44.48 57.80 61.83 63.98
GNN-PPI 63.81 74.72 71.37 82.67
ProtBert 70.94 73.36 70.32 78.86
ESM-1b 74.92 78.83 77.49 82.13
KeAP 78.58 77.54 77.22 $\textbf{84.74}^{*}$
GOProteinGNN $\textbf{79.14}^{*}$ $\textbf{80.31}^{*}$ $\textbf{78.11}^{*}$ 83.43
5.4. Semantic Similarity Inference
Table 4 presents the results of the semantic similarity inference task. In particular, our model outperformed all baselines significantly in capturing semantic similarities within the biological process (BP) category.
Biological processes play a crucial role in many drug development processes, making this category particularly important as it pertains directly to drug development. The success in this task underscores GOProteinGNN’s efficacy in acquiring enriched representations that encompass the complex biological processes in which proteins are involved, due to the whole protein knowledge graph learning, surpassing the capabilities of alternative approaches.
Furthermore, KeAP (Zhou et al., 2023) showed inferior performance compared to ESM-1b (Rives et al., 2019), which does not use factual knowledge. This result is surprising given that both GOProteinGNN and KeAP utilize the same factual knowledge repository, yet KeAP’s method of knowledge injection failed to outperform the model that does not use knowledge at all.
Table 4. Results on semantic similarity inference task. The best result is highlighted in bold.
Method Biological Process
MSA Transformer 0.31
ProtT5-XL 0.21
ProtBert 0.35
ESM-1b 0.42
KeAP 0.41
GOProteinGNN $\textbf{0.52}^{*}$
5.5. Protein–Protein Binding Affinity Estimation
Table 5 presents the results of the binding affinity estimation task. Evidently, GOProteinGNN outperformed PIPR (Chen et al., 2019) and ProtBert (Elnaggar et al., 2022) with a significant margin of improvement. Furthermore, it is noteworthy that GOProteinGNN exhibited a slight improvement over KeAP (Zhou et al., 2023).
The performances of ESM-1b (Rives et al., 2019) and GOProteinGNN are comparable, with no statistically significant differences observed between them. This similarity in performance may be attributed to the task’s sensitivity to subtle structural changes, which ESM-1b captures effectively through its extensive pre-training. In this context, the external factual knowledge leveraged by GOProteinGNN does not provide an advantage.
Table 5. Results on protein-protein binding affinity prediction task. Affinity (↓) means the lower the better. The best result is highlighted in bold.
@@ 5.5. Protein–Protein Binding Affinity Estimation
6. Ablation Tests
6.1. GNN Knowledge Injection (GKI) Ablation
The ablation test for the GKI component is presented in Table 6. The full model results highlight the effectiveness of GOProteinGNN’s model components, where the model demonstrated superior performance across all metrics, underscoring its pivotal role.
Ignoring relations through the GKI component implies using GCN (Kipf and Welling, 2017) instead of RGCN (Schlichtkrull et al., 2017) and ignoring edge type during the knowledge graph learning. Without relation learning, we observe a slight precision decrease, emphasizing its importance. Moreover, excluding the GKI component (i.e., without GNN) substantially degrades performance, reinforcing the significance of the GNN mechanism for effective knowledge integration.
These experiments highlight the contribution of the GKI component and the GNN mechanism to GOProteinGNN’s ability to capture essential biological contexts and interactions, leading to performance improvements.
Table 6. Ablation test for the GKI component on amino acid contact prediction. The metrics are precision scores (P@L, P@L/2, and P@L/5) for short- $(6\leq seq<12)$ contact predictions. The best result is highlighted in bold.
Method P@L P@L/2 P@L/5
GOProteinGNN $\textbf{0.43}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.65}^{*}$
-- GNN Ignoring Relations 0.40 0.51 0.64
-- Without GNN 0.30 0.40 0.52
6.2. Number of GKI Components Ablation
We wish to experiment with the number of optimal GKI components. Each GKI component adds an additional knowledge graph injection to the model. In the case of GOProteinGNN, the number of GKI components (referred to as $k$) is set to 3. In this ablation test, we reduce the number of GKI components to observe their impact. As shown in Table 7, reducing knowledge injections in GOProteinGNN results in a slight performance decrease. This decrease highlights the significance and meaningfulness of the GKI component.
Table 7. Ablation test for the number of GKI components ($k$) on amino acid contact prediction. The metrics are precision scores (P@L, P@L/2, and P@L/5) for short- $(6\leq seq<12)$ contact predictions. The best result is highlighted in bold.
Method P@L P@L/2 P@L/5
GOProteinGNN ($k=3$) $\textbf{0.43}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.65}^{*}$
GOProteinGNN ($k=2$) 0.42 0.53 0.63
7. Conclusions
In this paper, we introduced GOProteinGNN, a novel architecture that enhances PLMs by integrating protein knowledge graph information during the creation of amino acid-level representations. Specifically, GOProteinGNN leverages a novel GKI procedure that leverages the [CLS] token representation of the PLM during the encoder stage. Our methodology enables the incorporation of information at both the individual amino acid and protein levels, fostering a comprehensive learning process through graph-based techniques. Furthermore, GOProteinGNN uniquely learns the entire protein knowledge graph during training, allowing it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work.
GOProteinGNN showcases its potential across diverse protein-related tasks, such as amino acid contact prediction and protein-protein interaction identification. We posit that the observed performance improvements arise from a deeper comprehension of proteins themselves and the associated factual knowledge. This understanding allows generalization for never-before-seen proteins, as demonstrated on downstream tasks. The ability of GOProteinGNN to jointly model sequential data and knowledge graph information empowers it to handle complex biological scenarios effectively. This fusion of language model-based representation and knowledge graph processing makes GOProteinGNN a valuable tool for advancing research in the field of bioinformatics and drug development.
GOProteinGNN can enhance multiple real-world tasks, including protein structure prediction and drug-target interaction. By leveraging both the protein knowledge graph and sequential data modeling, GOProteinGNN effectively propagates information regarding proteins with biological connections. In drug development, it can facilitate virtual screening by prioritizing compounds for binding to specific protein targets, underscoring its potential to provide deeper insights into complex biological phenomena and expediting the drug discovery process.
For future directions, we envision the possibility of incorporating additional factual knowledge beyond proteins and extending our graph learning framework to more advanced methodologies. Additionally, exploring the applicability of our algorithm to general Natural Language Processing (NLP) tasks could be an intriguing avenue to explore.

### v2

@@ D.0.1. Contact Prediction
D.0.2. Semantic Similarity Inference
This task aims to evaluate how much information representation models capture about biomolecular functional similarity, with a specific emphasis on the biological process (BP) category within the Gene Ontology (GO) dataset. We adopt the semantic similarity measures outlined by Unsal et al. and compute Spearman’s rank correlation between representation vector similarities and the actual GO-based semantic similarities of corresponding protein pairs. Higher correlation values indicate greater success in representing functional similarity.
D.0.3. Protein-Protein Interaction Identification
Protein-protein interactions (PPI) refer to the physical contacts between two or more amino acid sequences, or two proteins, in our study. The goal is to predict the interaction type(s) of each protein pair, out of 7 possible interactions (i.e., multilabel classification task). The evaluation is conducted on the datasets: SHS27K (Chen et al., 2019) and SHS148K (Chen et al., 2019). Breadth-first search (BFS) is used to generate test sets from the two datasets. F1 score is used as the evaluation metric.
D.0.4. Remote Homology Detection
@@
Appendix E Protein-Protein Binding Affinity Estimation
In this section, we present a task where GOProteinGNN provides limited advantages over existing methods. The Protein-Protein Binding Affinity Estimation task involves evaluating the capability of a protein representation to predict changes in binding affinity resulting from protein mutations. In this context, GOProteinGNN does not incorporate protein structural information and thus offers no additional benefits compared to other methods. The results indicate that while performance is not degraded, it also does not show significant improvement.
E.0.1. Task Definition
@@
Each protein pair is associated with a real value in this task, making it a regression task that reflects binding affinity changes. We follow the approach of Unsal et al. and use the SKEMPI dataset (Moal and Fernández-Recio, 2012) with an evaluation based on the mean squared error obtained by 10-fold cross-validation.
E.0.2. Task Results
Table 7 presents the binding affinity estimation task results. GOProteinGNN outperformed PIPR (Chen et al., 2019) and ProtBert (Elnaggar et al., 2022) with a significant margin of improvement. Furthermore, it is noteworthy that GOProteinGNN exhibited a slight improvement over KeAP (Zhou et al., 2023). The performances of ESM-1b (Rives et al., 2019), ESM-2 (Lin et al., 2023), and GOProteinGNN are comparable, with no statistically significant differences. This similarity in performance may be attributed to the task’s sensitivity to subtle structural changes, which ESM-1b and ESM-2 capture effectively through their extensive pre-training, which are critical for this task (Unsal et al., 2022). In this context, the external factual knowledge leveraged by GOProteinGNN does not provide an advantage. However, the protein knowledge graph includes descriptions of structural properties, allowing GOProteinGNN to inject such knowledge directly into protein representations. This capability contributes to its improved understanding of the protein knowledge graph, leading to modest improvements over KeAP (Zhou et al., 2023).
@@ E.0.2. Task Results
Table 7. Results on protein-protein binding affinity prediction task. Affinity (↓) means the lower the better.